OpenStudy (anonymous):

Use the Change of Base Formula to evaluate log7 (76)

OpenStudy (anonymous):

\[\log_7(76)=\frac{\ln(76)}{\ln(7)}\] by the change of base formula

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

yw